Sec square(tan inverse 2) +cosec square(tan inverse 3) =

Grade:12th pass

1 Answers

Aryan Prajapati
7 years ago
Sec square(tan inverse 2) +cosec square(tan inverse 3)= 1+( tan (tan inverse 3))^2 + 1+ cot square(tan inverse 3) = 1+( tan (tan inverse 3))^2 + 1+ cot square( cot inverse 1/3)= 1+ 2*2 + 1 +(1/3)*(1/3)= 55/9